OpenStudy (anonymous):

he height in meters of a projectile can be modeled by h = −4.9t2 + vt + s where t is the time (in seconds) the object has been in the air, v is the initial velocity (in meters per seconds) and s is the initial height (in meters). A soccer ball is kicked upward from the ground and flies through the air with an initial vertical velocity of 4.9 meters per second. After how many seconds does it land approximately?

OpenStudy (anonymous):

*The

OpenStudy (anonymous):

You want the height to be 0 for the time it takes the ball to land (0 metres off ground). The equation will be 0=-4.9t^2 + 4.9t s=0 because the ball is kicked from the ground (0 metres off ground) if you divide both sides by -4.9 you get 0=(t^2) - t ===> this factorizes to t(t-1)=0 therefore t-1=0 therefore t=1 , it lands after 1 second